1) Calculate the change in entropy that occurs in the system when 4.50 mole of isopropyl alcohol (C3H8O) melts at its melting point (-89.5∘C). ΔH∘fus = 5.37 kJ/mol

2) Calculate the change in entropy that occurs in the system when 4.30 mole of diethyl ether (C4H6O) condenses from a gas to a liquid at its normal boiling point (34.6∘C). ΔHvap = 26.5 kJ/mol
1
Expert's answer
2020-04-15T01:33:02-0400

1)

"\Delta S =\frac{Q_{total}}{T}=\frac{\Delta H\times n }{T}=\frac{5.37\times10^3\times4.5}{273-89.5}=\frac{24.17\times10^3}{183.5}=131.72JK^{-1}"


2)

"\Delta S =\frac{Q_{total}}{T}=\frac{\Delta H\times n }{T}=\frac{26.5\times10^3\times4.3}{273+34.6}=\frac{113.95\times10^3}{317.6}=358.78JK^{-1}"